# To Make, or to Take, That Is the Question: Impact of LOB Mechanics on Natural Trading Strategies

### v1

@@ 6 Fill and Be Killed?
Figure 11 demonstrates the relationship. The horizontal axis labels the on-paper returns of all the orders in our sample (whether filled or canceled because the top price moved away); returns are measured from the time of the order’s terminal state (fill-or-cancel) five seconds into the future. The vertical axis shows the proportion of those orders that filled; that is, an estimate of the fill probability given the subsequent return. The plot reveals a notable anti-correlation: orders with negative subsequent five-second returns are highly likely to fill, whereas those with positive returns are much less likely to fill. While this result alone does not guarantee an unconditional expectation of a negative return—since the negative-return orders might represent only a small portion of the entire sample, with their large negative returns being counterbalanced by small positive returns from a larger number of orders—our empirical data does show a negative expected return of approximately $-0.8$ bp, or $-0.3$ bp net of rebate.
Figure 11: Empirical relationship between fill probability and return measured five-second from post-time.
@@
A common misconception vitiating the HFT and market making literature is the belief that market makers can simply post orders at the top bid and ask prices to earn the spread plus rebates. In a number of such papers (see for instance Cartea [2016], Avellaneda and Stoikov [2008]), it is commonly assumed that those maker orders fill randomly at Poisson rates, with no impact on the subsequent mid-price evolution, which is often modeled as a continuous diffusion process.
@@
The cruel reality is that such market making strategies are a quick path to bankruptcy due to negative price drift, which we now illustrate with a case study of the prototypical naive market making strategy. We mirror the approach of DeLise [2024], who carried out a similar analysis of the same strategy in US treasury markets, and it is interesting to note that the same results hold in our completely different asset class. That strategy can be specified as follows:
@@
Basic MM Strategy Step 1: Post a minimum-sized buy order at the top bid and a minimum-sized sell order at the top ask. Step 2: If the position differs from zero (because one of the two initial orders has filled), post only on the side that will return the position to zero: Keep the unfilled order open while it is posted at the touch. Should the price change without the order being filled, so it no longer being posted at the top price, cancel the order and place a new one at the new top price. Step 3: Once a fill is received that returns the inventory back to zero, return to Step 1.
@@
One might imagine that by buying low and selling higher (by the spread), the strategy would capture the spread and earn twice the maker rebate per pair of posts. However, this is only true if both orders are filled without any change in the midprice. In reality, it is far more common that, after one order has filled, an adverse midprice movement forces the cancellation and reposting of the unfilled order at a worse price. We use our empirical data to assess the impact of this fact.
@@
In our overall experiment, we continuously quoted minimum-sized orders on both bid and ask without inventory restrictions. The orders submitted by the basic market-making strategy are a subset of the whole. We extract these and analyze the performance of the strategy.111111 Our results with minimum-sized orders are indicative of what one might expect with larger order sizes. One might reasonably expect results to deteriorate with larger order sizes: the literature on market impact tells us that the market impact of taker orders grows with their size. Hence, since the market impact of a taker order is the same as that of adverse selection on the maker order on the other end of the trade, the subsequent expected return of a maker order ought to decrease in its filled amount.
@@
In Figure 13 we show the strategy’s cumulative return (including maker rebates) in basis points over a period of approximately three days and seven hours. By the end of this period, the strategy has lost almost 60%; in annualized units, the Sharpe ratio is a staggering $-109.0$. The average holding time between opening and closing of a position is approximately 11 seconds.
@@
Figure 13 illustrates the distribution of realized spreads (the price difference in basis points between an opening fill that changes the position from 0 to ±$\pm 1$, and the subsequent rebalancing fill) with the frequency on a logarithmic scale. While the most frequent outcome for a round-trip is indeed a modest positive return, the strategy is plagued by a high occurrence of larger losses, and occasional catastrophic ones, which overwhelm the small gains.
@@
The canonical example of a trading signal, widely accepted in the microstructure literature, is the order book imbalance; indeed, we have referred to it repeatedly above. We investigate a set of imbalance-based strategies commonly found in the literature (see for instance Cartea et al. [2018], Cont et al. [2023]) which, while still relatively naive, also have prima facie appeal. We show that this, too, is a false hope.
@@
  • Imbalance-based maker strategy: Post maker orders in the direction of the imbalance. This strategy differs from the naive one by only submitting new orders when the imbalance is favorable to the maker order (Imbalance threshold to post: magnitude $>0.5$).
  • Imbalance-based maker strategy with cancellation: Similar to the above, but with an added rule to cancel posted orders if the imbalance shifts adversely to the order (Imbalance threshold to post: magnitude $>0.5$; cancel threshold: 0.0).
  • Imbalance-based taker strategy: Trade in the direction of the imbalance using taker orders, employing the same inventory-balancing approach as the maker strategies (Imbalance threshold to submit a taker order: magnitude $>0.5$).  [^12] Latency is ignored here, potentially leading to an overstatement of results, see Albers et al. [2024].
@@
Imbalance taker 18927 $-1.9621$ 2.1653 15.04
Naive maker 12623 $-0.4307$ 2.6486 22.56
Imbalance maker w/o cancel 8851 $-0.4705$ 2.8529 32.17
Imbalance maker w/ cancel 7333 $-0.4921$ 3.1143 38.83
@@ 7 Reversals
We noted in Section 4 that orders posted when the opposite-side queue is large, and the near-side queue is small (implying an adverse imbalance) are associated high fill probabilities. While those conditions usually imply a poor expected return, as established in Section 5, there is a subset of cases where the initially adverse imbalance shifts favorably after post-time, i.e. when the initial imbalances falsely predicts the next price change. These scenarios thus offer the dual benefits of high fill probability and positive subsequent return, rendering them especially significant for market makers - a class of scenarios we denote as reversals.
Figure 16 illustrates a confusion matrix of the imbalance as a predictor of the next price change. Orders trading against the imbalance in the off-diagonal cases (reversals), where the imbalance is a false predictor, have a high fill probability and yield a favorable subsequent return.
@@ 7.1 Modelling Approach
    The order achieves a positive short-term return: let $T^{\prime}\geq T$ denote the timestamp of the first post-fill price change. The requirement is: $\operatorname{ret}_{T,T^{\prime}}(\mathfrak{o}):=\operatorname{sgn}(\mathfrak{o})\left(p(\mathfrak{o})/p_{T^{\prime}}-1\right)\cdot 10,000>0$. Here, $p(\mathfrak{o})$ denotes the order’s limit price, while $p_{T^{\prime}}$ is the top bid price at time $T^{\prime}$ for a buy order or the top ask price for a sell order, with the signed return accounting for order direction using $\operatorname{sgn}(\mathfrak{o})=1$ for buys and $-1$ for sells.
@@ 7.2 Features
In this subsection, we outline the features we use to predict reversals. We expect this to be inherently challenging due to the ease with which accurate predictions can be parlayed into trading profits (cf. Unprofitability Principle, Section 1.1.2). Although the literature has not directly addressed our specific problem, we can draw insights from related prediction tasks, such as short-term return predictions. For instance, Ait-Sahalia et al. [2022] constructs non-overlapping lookback windows to calculate various variables. Similarly, Cartea et al. [2023] develops a feature set designed to capture the current state of the LOB and recent LOB activity, which they use to forecast market participants’ trading decisions.
@@
Furthermore, our features are pre-processed such that the model need not differentiate between the direction of the order (buy or sell). Concretely this means that instead of using "buy" and "sell" we use "same-direction" and "opposite-direction", or instead of "bid" and "ask" we use "near-side" and "opposite-side", and for some other features such as returns it means multiplying them by a sign based on the direction of the order. All of this allows the model to treat buys and sells equivalently during training and prediction.
@@ 7.3 Results
    Balanced-Inventory Strategy: This strategy operates similarly to the basic MM strategy described in Section 6, but it restricts order submissions to situations where the model (logistic regression or random forest) predicts a reversal probability exceeding a threshold $p\in[0,1]$. Specifically, it posts a buy order at the top bid price if a (downward) reversal is predicted with probability $>p$, and a sell order at the top ask price if an (upward) reversal is predicted with probability $>p$. If the inventory is non-zero (i.e., long or short one unit), the strategy only posts on the side that returns the position to zero, provided a reversal is predicted with probability $>p$. Since this strategy maintains a balanced inventory (similar to the basic MM strategy), we can meaningfully compute its realized PnL (unlike the Undirectional Strategy).
@@
Threshold Logistic Regression Random Forest # Orders / Day Fill Prob Return (5s) Stdev p-value # Orders / Day Fill Prob Return (5s) Stdev p-value 0.00 21689 0.8824 $-0.74$ 1.55 0.5000 21657 0.8824 $-0.74$ 1.55 0.5000 0.02 19438 0.8809 $-0.71$ 1.62 0.0003 21402 0.8825 $-0.71$ 1.56 0.0001 0.04 16596 0.8791 $-0.72$ 1.72 0.0165 20780 0.8813 $-0.67$ 1.57 0.0000 0.06 13939 0.8773 $-0.71$ 1.82 0.0018 19733 0.8807 $-0.67$ 1.58 0.0000 0.08 11674 0.8767 $-0.72$ 1.92 0.0397 18375 0.8794 $-0.68$ 1.61 0.0000 0.10 9566 0.8770 $-0.72$ 2.05 0.0607 16633 0.8785 $-0.68$ 1.66 0.0000 0.12 7723 0.8778 $-0.70$ 2.17 0.0040 14790 0.8773 $-0.69$ 1.71 0.0000 0.14 5979 0.8829 $-0.70$ 2.28 0.0189 12848 0.8761 $-0.71$ 1.77 0.0008 0.16 4441 0.8855 $-0.69$ 2.43 0.0083 11006 0.8743 $-0.71$ 1.83 0.0034 0.18 3158 0.8880 $-0.69$ 2.57 0.0334 9304 0.8732 $-0.69$ 1.90 0.0000 0.20 2104 0.8887 $-0.68$ 2.77 0.0496 7732 0.8690 $-0.70$ 1.96 0.0017 0.22 1365 0.8814 $-0.68$ 3.00 0.0903 6265 0.8687 $-0.69$ 2.04 0.0006 0.24 882 0.8673 $-0.62$ 3.24 0.0278 5034 0.8642 $-0.68$ 2.13 0.0003 0.26 577 0.8612 $-0.60$ 3.55 0.0432 3965 0.8589 $-0.68$ 2.18 0.0018 0.28 349 0.8559 $-0.45$ 3.86 0.0055 3037 0.8572 $-0.67$ 2.27 0.0010 0.30 233 0.8545 $-0.32$ 3.78 0.0011 2271 0.8523 $-0.63$ 2.41 0.0000 0.32 162 0.8284 $-0.40$ 3.78 0.0187 1675 0.8548 $-0.57$ 2.52 0.0000 0.34 113 0.8128 $-0.27$ 3.82 0.0093 1219 0.8527 $-0.50$ 2.59 0.0000 0.36 69 0.7913 $+0.27$ 3.49 0.0000 837 0.8507 $-0.41$ 2.61 0.0000 0.38 45 0.7733 $+1.34$ 3.57 0.0000 582 0.8521 $-0.32$ 2.80 0.0000 0.40 33 0.7818 $+2.11$ 3.75 0.0000 375 0.8465 $-0.33$ 2.99 0.0000 0.42 24 0.8250 $+2.65$ 3.69 0.0000 248 0.8537 $-0.30$ 3.12 0.0000 0.44 19 0.8125 $+2.66$ 3.72 0.0000 159 0.8441 $-0.13$ 3.54 0.0000 0.46 13 0.7826 $+3.45$ 3.88 0.0000 96 0.8113 $+0.05$ 2.88 0.0000 0.48 11 0.8421 $+3.49$ 4.23 0.0000 52 0.8161 $+0.76$ 3.02 0.0000 0.50 9 0.8750 $+3.02$ 1.92 0.0000 23 0.7692 $+1.68$ 3.35 0.0000
How can we ascertain if the improvements in post-fill returns shown in Table 3 genuinely surpass the performance of randomly placed orders, considering the possibility that these results might be mere coincidences? This can be addressed through a two-sample t-test, which compares the average returns of two groups: the strategy for a given model and threshold versus random orders. For instance, at the $0.30$ reversal probability threshold for the Logistic Regression model, where the average post-fill return is $-0.32$, we test the null hypothesis that there is no difference between the returns of this trading strategy and random orders. The calculated t-statistic is $3.062602$ with a p-value of $0.0011$, allowing us to reject the null hypothesis and conclude that the observed return improvement is statistically significant at the 5% level. The table also reports p-values for all reversal probability thresholds, providing a view of the statistical significance across different levels.
@@
Let us now examine Table 4, which presents the performance of the Balanced Inventory strategy based on a model and a specified reversal probability threshold. At a threshold of 00, the strategy operates without restrictions, and is equivalent to the naive market-making strategy discussed in Section 6. As the threshold increases, the strategy restricts order submissions to scenarios where the model predicts a reversal with a certain probability, resulting in fewer orders being posted.
The returns are calculated as realized roundtrip returns, defined as the price difference (in bp) between the limit prices of two orders: the initial order that accumulates inventory (changing it from 0 to ±$\pm 1$) and the subsequent order that returns the inventory to 0.151515Returns are signed such that a positive return indicates a profit, while a negative return indicates a loss. Each leg of the roundtrip earns a maker rebate of $0.5$ bp, thus we add $1$ bp to each realized roundtrip return. We also present the standard deviations of these returns, the average holding time (the duration between an inventory-acquiring trade and its subsequent hedge), and the annualized Sharpe ratio of the strategy. The Sharpe ratio is calculated based on hourly periods, given that our test period spans approximately 3.5 trading days.
@@
The improvement in realized roundtrip PnL compared with the naive strategy is striking. The naive strategy incurs an average loss of 0.44 basis points per roundtrip (net of rebates), occurring at a high frequency, which explains its poor Sharpe ratio. At a higher threshold, such as 0.24 for the Logistic Regression model, the number of roundtrips remains relatively large (327 per day), and the average return (0.71) constitutes a significant improvement over the naive strategy. A hypothesis test similar to the one conducted earlier confirms that this difference is statistically significant and unlikely to be due to chance, despite the large standard deviation of returns. The cumulative PnL plot and return histogram for this threshold are shown in Figures 18 and 18, respectively, illustrating the strategy’s improved performance and the distribution of realized returns; these are the analogues of the plots for the naive MM strategy shown in Figures 13 and 13. The Sharpe ratio improves from $-109$ for the naive strategy to 11.97 for the Logistic Regression-based strategy at the 0.24 threshold, demonstrating unequivocally that our model yields substantial economic benefit. It is important to note that the impressive Sharpe ratio is unlikely to persist with larger-than-minimum-sized orders, and that our calculations assume the best possible maker rebate (0.5 bp). Even minor changes to these assumptions could result in a significantly lower Sharpe ratio, yet it remains far superior to that of the naive strategy.

### v2

@@ 6  Fill and Be Killed?
Figure 11 demonstrates the relationship. The horizontal axis labels the on-paper returns of all the orders in our sample (whether filled or canceled because the top price moved away); returns are measured from the time of the order’s terminal state (fill-or-cancel) five seconds into the future. The vertical axis shows the proportion of those orders that filled; that is, an estimate of the fill probability given the subsequent return. The plot reveals a notable anti-correlation: orders with negative subsequent five-second returns are highly likely to fill, whereas those with positive returns are much less likely to fill. While this result alone does not guarantee an unconditional expectation of a negative return—since the negative-return orders might represent only a small portion of the entire sample, with their large negative returns being counterbalanced by small positive returns from a larger number of orders—our empirical data does show a negative expected return of approximately 0.8-0.8 bp, or 0.3-0.3 bp net of rebate.
Figure 11: Empirical relationship between fill probability and return measured five seconds from post-time.
@@
A common misconception vitiating the HFT and market making literature is the belief that market makers can simply post orders at the top bid and ask prices to earn the spread plus rebates. In a number of such papers (see for instance Cartea_Jaimungal_2016; Avellaneda_2008), it is commonly assumed that those maker orders fill randomly at Poisson rates, with no impact on the subsequent mid-price evolution, which is often modeled as a continuous diffusion process.
@@
The cruel reality is that such market making strategies are a quick path to bankruptcy due to negative price drift, which we now illustrate with a case study of the prototypical naive market making strategy. We mirror the approach of DeLise:2024aa, who carried out a similar analysis of the same strategy in US treasury markets, and it is interesting to note that the same results hold in our completely different asset class. That strategy can be specified as follows:
@@
Basic $M$ Strategy Step 1: Post a minimum-sized buy order at the top bid and a minimum-sized sell order at the top ask. Step 2: If the position differs from zero (because one of the two initial orders has filled), post only on the side that will return the position to zero: Keep the unfilled order open while it is posted at the touch. Should the price change without the order being filled, so it no longer being posted at the top price, cancel the order and place a new one at the new top price. Step 3: Once a fill is received that returns the inventory back to zero, return to Step 1.
@@
One might imagine that by buying low and selling higher (by the spread), the strategy would capture the spread and earn twice the maker rebate per pair of posts. However, this is only true if both orders are filled without any change in the midprice. The strategy described above is ”naive” primarily because it lacks a sophisticated cancellation policy; orders are left in the book even when a skilled trader might cancel them to avoid adverse selection. In reality, it is far more common that, after one order has filled, an adverse midprice movement forces the cancellation and reposting of the unfilled order at a worse price. We use our empirical data to assess the impact of this fact and later illustrate how a simple, signal-based cancellation policy can mitigate, but not eliminate, these losses.
@@
In our overall experiment, we continuously quoted minimum-sized orders on both bid and ask without inventory restrictions. The orders submitted by the basic market-making strategy are a subset of the whole. We extract these and analyze the performance of the strategy. [^13] Our results with minimum-sized orders are indicative of what one might expect with larger order sizes. One might reasonably expect results to deteriorate with larger order sizes: the literature on market impact tells us that the market impact of taker orders grows with their size. Hence, since the market impact of a taker order is the same as that of adverse selection on the maker order on the other end of the trade, the subsequent expected return of a maker order ought to decrease in its filled amount.
@@
In Figure 13 we show the strategy’s cumulative return (including maker rebates) in basis points over a period of approximately three days and seven hours. By the end of this period, the strategy has lost almost 60%; in annualized units, the Sharpe ratio is a staggering 109.0-109.0. The average holding time between opening and closing of a position is approximately 11 seconds.
@@
Figure 13 illustrates the distribution of realized spreads (the price difference in basis points between an opening fill that changes the position from 0 to $\pm 1$, and the subsequent rebalancing fill) with the frequency on a logarithmic scale. While the most frequent outcome for a round-trip is indeed a modest positive return, the strategy is plagued by a high occurrence of larger losses, and occasional catastrophic ones, which overwhelm the small gains.
@@
A canonical example of a trading signal, widely accepted in the microstructure literature, is the order book imbalance; indeed, we have referred to it repeatedly above. We investigate a set of imbalance-based strategies commonly found in the literature (see for instance Cartea_2018aa; Cont_2023) which, while still relatively naive, also have prima facie appeal. We show that this, too, is a false hope.
@@
  • Imbalance-based maker strategy: Post maker orders in the direction of the imbalance. This strategy differs from the naive one by only submitting new orders when the imbalance is favorable to the maker order (Imbalance threshold to post: magnitude >0.5>0.5).
  • Imbalance-based maker strategy with cancellation: Similar to the above, but with an added rule to cancel posted orders if the imbalance shifts adversely to the order (Imbalance threshold to post: magnitude >0.5>0.5; cancel threshold: 0.0).
  • Imbalance-based taker strategy: Trade in the direction of the imbalance using taker orders, employing the same inventory-balancing approach as the maker strategies (Imbalance threshold to submit a taker order: magnitude >0.5>0.5).  [^14] Latency is ignored here, potentially leading to an overstatement of results, see Albers_2024.
@@
Imbalance taker 18927 1.9621-1.9621 2.1653 15.04
Naive maker 12623 0.4307-0.4307 2.6486 22.56
Imbalance maker w/o cancel 8851 0.4705-0.4705 2.8529 32.17
Imbalance maker w/ cancel 7333 0.4921-0.4921 3.1143 38.83
@@ 7  Reversals
We noted in Section 4 that orders posted when the opposite-side queue is large, and the near-side queue is small (implying an adverse imbalance) are associated high fill probabilities. While those conditions usually imply a poor expected return, as established in Section 5, there is a subset of cases where the initially adverse imbalance shifts favorably after post-time, i.e. when the initial imbalances falsely predicts the next price change. These scenarios, which we term reversals, represent instances where the typical negative correlation between fill probability and post-fill returns can break down. While not a fully-fledged trading strategy, identifying such reversals offers a potential resolution to the fundamental challenge facing maker orders. We present one such model not as a definitive solution, but to illustrate the mechanics involved and demonstrate that such opportunities can, in principle, be identified.
Figure 16 shows a confusion matrix of the imbalance as a predictor of the next price change. Orders trading against the imbalance in the off-diagonal cases (reversals), where the imbalance is a false predictor, have a high fill probability and yield a favorable subsequent return.
@@ 7.1  Modelling Approach
    The order achieves a positive short-term return: let $T^{\prime}\geq T$ denote the timestamp of the first post-fill price change. The requirement is: $\operatorname{ret}_{T,T^{\prime}}(\mathfrak{o}):=\operatorname{sgn}(\mathfrak{o})\left(p(\mathfrak{o})/p_{T^{\prime}}-1\right)\cdot 10,000>0$. Here, $p(\mathfrak{o})$ denotes the order’s limit price, while $p_{T^{\prime}}$ is the top bid price at time $T^{\prime}$ for a buy order or the top ask price for a sell order, with the signed return accounting for order direction using $\operatorname{sgn}(\mathfrak{o})=1$ for buys and 1-1 for sells.
@@ 7.2  Features
In this subsection, we outline the features we use to predict reversals. We expect this to be inherently challenging due to the ease with which accurate predictions can be parlayed into trading profits (cf. Unprofitability Principle, Section 1.1.2). Although the literature has not directly addressed our specific problem, we can draw insights from related prediction tasks, such as short-term return predictions. For instance, Ait_Sahalia_2022 constructs non-overlapping lookback windows to calculate various variables. Similarly, Cohen_2023 develops a feature set designed to capture the current state of the LOB and recent LOB activity, which they use to forecast market participants’ trading decisions.
@@
Furthermore, our features are pre-processed such that the model need not differentiate between the direction of the order (buy or sell). Concretely this means that instead of using ”buy” and ”sell” we use ”same-direction” and ”opposite-direction”, or instead of ”bid” and ”ask” we use ”near-side” and ”opposite-side”, and for some other features such as returns it means multiplying them by a sign based on the direction of the order. All of this allows the model to treat buys and sells equivalently during training and prediction.
@@ 7.3  Results
    Balanced-Inventory Strategy: This strategy operates similarly to the basic MM strategy described in Section 6, but it restricts order submissions to situations where the model (logistic regression or random forest) predicts a reversal probability exceeding a threshold $p\in[0,1]$. Specifically, it posts a buy order at the top bid price if a (downward) reversal is predicted with probability >p>p, and a sell order at the top ask price if an (upward) reversal is predicted with probability >p>p. If the inventory is non-zero (i.e., long or short one unit), the strategy only posts on the side that returns the position to zero, provided a reversal is predicted with probability >p>p. Since this strategy maintains a balanced inventory (similar to the basic MM strategy), we can meaningfully compute its realized PnL (unlike the Undirectional Strategy).
@@
Threshold Logistic Regression Random Forest # Orders / Day Fill Prob Return (5s) Stdev p-value # Orders / Day Fill Prob Return (5s) Stdev p-value 0.00 21689 0.8824 0.74-0.74 1.55 0.5000 21657 0.8824 0.74-0.74 1.55 0.5000 0.02 19438 0.8809 0.71-0.71 1.62 0.0003 21402 0.8825 0.71-0.71 1.56 0.0001 0.04 16596 0.8791 0.72-0.72 1.72 0.0165 20780 0.8813 0.67-0.67 1.57 0.0000 0.06 13939 0.8773 0.71-0.71 1.82 0.0018 19733 0.8807 0.67-0.67 1.58 0.0000 0.08 11674 0.8767 0.72-0.72 1.92 0.0397 18375 0.8794 0.68-0.68 1.61 0.0000 0.10 9566 0.8770 0.72-0.72 2.05 0.0607 16633 0.8785 0.68-0.68 1.66 0.0000 0.12 7723 0.8778 0.70-0.70 2.17 0.0040 14790 0.8773 0.69-0.69 1.71 0.0000 0.14 5979 0.8829 0.70-0.70 2.28 0.0189 12848 0.8761 0.71-0.71 1.77 0.0008 0.16 4441 0.8855 0.69-0.69 2.43 0.0083 11006 0.8743 0.71-0.71 1.83 0.0034 0.18 3158 0.8880 0.69-0.69 2.57 0.0334 9304 0.8732 0.69-0.69 1.90 0.0000 0.20 2104 0.8887 0.68-0.68 2.77 0.0496 7732 0.8690 0.70-0.70 1.96 0.0017 0.22 1365 0.8814 0.68-0.68 3.00 0.0903 6265 0.8687 0.69-0.69 2.04 0.0006 0.24 882 0.8673 0.62-0.62 3.24 0.0278 5034 0.8642 0.68-0.68 2.13 0.0003 0.26 577 0.8612 0.60-0.60 3.55 0.0432 3965 0.8589 0.68-0.68 2.18 0.0018 0.28 349 0.8559 0.45-0.45 3.86 0.0055 3037 0.8572 0.67-0.67 2.27 0.0010 0.30 233 0.8545 0.32-0.32 3.78 0.0011 2271 0.8523 0.63-0.63 2.41 0.0000 0.32 162 0.8284 0.40-0.40 3.78 0.0187 1675 0.8548 0.57-0.57 2.52 0.0000 0.34 113 0.8128 0.27-0.27 3.82 0.0093 1219 0.8527 0.50-0.50 2.59 0.0000 0.36 69 0.7913 +0.27+0.27 3.49 0.0000 837 0.8507 0.41-0.41 2.61 0.0000 0.38 45 0.7733 +1.34+1.34 3.57 0.0000 582 0.8521 0.32-0.32 2.80 0.0000 0.40 33 0.7818 +2.11+2.11 3.75 0.0000 375 0.8465 0.33-0.33 2.99 0.0000 0.42 24 0.8250 +2.65+2.65 3.69 0.0000 248 0.8537 0.30-0.30 3.12 0.0000 0.44 19 0.8125 +2.66+2.66 3.72 0.0000 159 0.8441 0.13-0.13 3.54 0.0000 0.46 13 0.7826 +3.45+3.45 3.88 0.0000 96 0.8113 +0.05+0.05 2.88 0.0000 0.48 11 0.8421 +3.49+3.49 4.23 0.0000 52 0.8161 +0.76+0.76 3.02 0.0000 0.50 9 0.8750 +3.02+3.02 1.92 0.0000 23 0.7692 +1.68+1.68 3.35 0.0000
How can we ascertain if the improvements in post-fill returns shown in Table 3 genuinely surpass the performance of randomly placed orders, considering the possibility that these results might be mere coincidences? This can be addressed through a two-sample t-test, which compares the average returns of two groups: the strategy for a given model and threshold versus random orders. For instance, at the $0.30$ reversal probability threshold for the Logistic Regression model, where the average post-fill return is 0.32-0.32, we test the null hypothesis that there is no difference between the returns of this trading strategy and random orders. The calculated t-statistic is $3.062602$ with a p-value of $0.0011$, allowing us to reject the null hypothesis and conclude that the observed return improvement is statistically significant at the 5% level. The table also reports p-values for all reversal probability thresholds, providing a view of the statistical significance across different levels.
@@
Let us now examine Table 4, which presents the performance of the Balanced Inventory strategy based on a model and a specified reversal probability threshold. At a threshold of 0, the strategy operates without restrictions, and is equivalent to the naive market-making strategy discussed in Section 6. As the threshold increases, the strategy restricts order submissions to scenarios where the model predicts a reversal with a certain probability, resulting in fewer orders being posted.
The returns are calculated as realized roundtrip returns, defined as the price difference (in bp) between the limit prices of two orders: the initial order that accumulates inventory (changing it from 0 to $\pm 1$) and the subsequent order that returns the inventory to 0.171717Returns are signed such that a positive return indicates a profit, while a negative return indicates a loss. Each leg of the roundtrip earns a maker rebate of $0.5$ bp, thus we add $1$ bp to each realized roundtrip return. We also present the standard deviations of these returns, the average holding time (the duration between an inventory-acquiring trade and its subsequent hedge), and the annualized Sharpe ratio of the strategy. The Sharpe ratio is calculated based on hourly periods, given that our test period spans approximately 3.5 trading days.
@@
The improvement in realized roundtrip PnL compared with the naive strategy is striking. The naive strategy incurs an average loss of 0.44 basis points per roundtrip (net of rebates), occurring at a high frequency, which explains its poor Sharpe ratio. At a higher threshold, such as 0.24 for the Logistic Regression model, the number of roundtrips remains relatively large (327 per day), and the average return (0.71) constitutes a significant improvement over the naive strategy. A hypothesis test similar to the one conducted earlier confirms that this difference is statistically significant and unlikely to be due to chance, despite the large standard deviation of returns. The cumulative PnL plot and return histogram for this threshold are shown in Figures 18 and 18, respectively, illustrating the strategy’s improved performance and the distribution of realized returns; these are the analogues of the plots for the naive MM strategy shown in Figures 13 and 13. The Sharpe ratio improves from 109-109 for the naive strategy to 11.97 for the Logistic Regression-based strategy at the 0.24 threshold, demonstrating unequivocally that our model yields substantial economic benefit. It is important to note that the impressive Sharpe ratio is unlikely to persist with larger-than-minimum-sized orders, and that our calculations assume the best possible maker rebate (0.5 bp). Even minor changes to these assumptions could result in a significantly lower Sharpe ratio, yet it remains far superior to that of the naive strategy.